# A Regret Perspective on Online Selective Generation

### v1

@@ B.1 Online Learning and EW Algorithm
[14, 13] Let $\ell_{t}(\cdot)\in[0,\ell_{\text{max}}]$. For any $T\in\mathbb{N}$ and finite hypotheses $\mathcal{{H}}$, Algorithm 2 provides the following regret bound if $\eta=\sqrt{{8\ln|\mathcal{{H}}|}/{(\ell_{\text{max}}^{2}T)}}$:
$\textbf{Reg}_{T}\leq\ell_{\text{max}}\sqrt{{T\ln|\mathcal{{H}}|}/{2}}=\mathcal{{O}}(\ell_{\text{max}}\sqrt{T\ln|\mathcal{{H}}|}).$ (9)
@@
Appendix C Additional Method
C.1 Online Selective Generation with Full Feedback
Here, we specifically leverage the conventional exponential weighting (EW) algorithm [14] for online learning under adaptive adversaries to re-purpose it for addressing online selective generation with full feedback. This is an oracle method with the performance upper bound of our method with partial feedback. To this end, we provide the reduction of online selective generation with full feedback to the online learning problem. Then, we introduce a modified EW algorithm for selective generation, called EW-SG, followed by its regret bound. Finally, we leverage the proposed conversion lemma, which converts the regret bound of any online learning algorithms into the FDR bound, to show the FDR controllability of EW-SG at a desired level.
C.1.1 Reduction: From Online Selective Generation to Online Learning
@@
models selective generators $\mathcal{{H}}$ hypotheses $\mathcal{{H}}$
@@
C.1.2 Algorithm and Its Regret Bound
@@ C.2 Regret-to-FDR Conversion
Appendix D Algorithms
Algorithm 2 Exponential Weighting (EW) [14, 13]
@@ Appendix D Algorithms
Algorithm 3 Exponential-weight algorithm for Exploration and Exploitation (Exp3) [7]
1:procedure Exp3($T,\mathcal{{H}},\eta$)
@@
7:         Update $w_{t+1}(\tau)\propto w_{t}(\tau)\cdot\exp(-\eta\hat{\ell}_{t}(\tau))$ for all $\tau\in\mathcal{{H}}$
@@
9:         $\ell_{t}\left(\tau,\alpha\right)\leftarrow a_{t}(\tau)+\lambda d_{t}(\tau,\alpha)$ for all $\tau\in\mathcal{{H}}$
@@
Algorithm 5 Exp3 for Online Selective Generation with Partial Feedback (Exp3-SG)
1:procedure Exp3-SG($T,\mathcal{{H}},\alpha,\lambda,\eta$)
@@
5:         Predict $\hat{{S}}_{t}(\mathbf{x}_{t};\tau_{t})$ where $\tau_{t}\sim p_{t}(\tau)=\sum_{\tau\in\mathcal{{H}}}\frac{w_{t}(\tau)}{\sum_{\tau\in\mathcal{{H}}}w_{t}(\tau)}\cdot\delta(\tau)$ ()arm $\left(\triangleright\right)\text{arm selection}$
@@
10:procedure ComputeLoss($\hat{\mathbf{s}},b,\alpha,\lambda$)
11:     $a\leftarrow\mathbbm{1}(\hat{\mathbf{s}}=\texttt{IDK})$
12:     $d\leftarrow\mathbbm{1}(\hat{\mathbf{s}}\neq\texttt{IDK})\cdot b-\alpha\,\mathbbm{1}(\hat{\mathbf{s}}\neq\texttt{IDK})+\alpha$
13:     $\ell\leftarrow a+\lambda d$
14:     return $\ell$
Appendix E Experiment Setup
E.1 Computing Environment
@@
E.2 Distribution-shift Environment Setup
@@
Gradual Shift. This setup generates a smooth distribution change over time by drawing individual samples probabilistically between NQ and TriviaQA, with the sampling probability gradually shifting from pure NQ (or TriviaQA) at the start to pure TriviaQA (or NQ) at the end. In particular, given a multinomial distribution $[\frac{t}{T},T-\frac{t}{T}]$ for each dataset that changes over time, we first randomly choose a dataset which follows this multinomial distribution, and then we sample one question-answering pair in the sampled dataset. Over the full sequence length, this creates linear interpolation between the two datasets, enabling analysis of model performance under a progressive distribution shift.
E.3 Interactive Environment Setup
@@
To instruct the user-acting agent to generate questions, we use the SQuAD [24] dataset, which consists of many question-answer pairs, each associated with a gold context. Here, we discard the QA pairs and only use $20,962$ context passages, which vary over a range of topics (e.g., science, history, celebrities, entertainment, etc.), forming a diverse set of natural distributions. During the simulation, we randomly select a context from the dataset and provide it to the user-acting agent, which is prompted to continuously generate questions based on the given context. (e.g., Figure 1). To mimic a dynamically shifting environment, e.g., the user’s interest shifts over time, we further instruct the user-acting agent to change the context during the interaction when it determines it is done with the current one (e.g., "When you are done with the current context, say ’SHIFT’ to request a new context.") or after at most 10 question-answering turns. To design this custom workflow, we implement the core functions of the simulation, including memory handling and multi-turn conversation, using LangGraph, a framework built on LangChain [25].
Appendix F Additional Experiments
F.1 Stochastic Environment
@@
Figure 8: Comparison of selective generation methods under a stochastic environment with GPT-3.5-turbo as a generator on NQ ($T=30\mathrm{K},{\alpha=0.25}$). The violin plots are drawn with randomly chosen $30\mathrm{K}$ samples over $30$ random trials.
@@
Figure 9: Comparison of selective generation methods under a stochastic environment with LLaMA3.1-8B-Instruct as a generator on NQ ($T=30\mathrm{K},{\alpha=0.25}$). The violin plots are drawn with randomly chosen $30\mathrm{K}$ samples over $30$ random trials. Comparison of selective generation methods under stochastic environments with varying feedback.
(a) FDR over time
(b) Inefficiency over time
Figure 10: Comparison of selective generation methods under a stochastic environment with GPT-3.5-turbo as a generator on NQ ($T=30\mathrm{K},{\alpha=0.25}$) and $f_{\texttt{std}}$ as our scoring function. The violin plots are drawn with randomly chosen $30\mathrm{K}$ samples over $30$ random trials.
(a) FDR over time
(b) Inefficiency over time
Figure 11: Comparison of selective generation methods under a stochastic environment with LLaMA3.1-8B-Instruct as a generator on NQ ($T=30\mathrm{K},{\alpha=0.25}$) and $f_{\texttt{std}}$ as our scoring function. The violin plots are drawn with randomly chosen $30\mathrm{K}$ samples over $30$ random trials.
F.2 Distribution-shift Environment
F.2.1 Single Shift
@@
Figure 12: Comparison of selective generation methods under a distribution-shift environment with GPT-3.5-turbo as a generator ($T=120\mathrm{K},\alpha=0.1$). Here, we consider a single distribution shift from $60\mathrm{K}$-sized NQ to $60\mathrm{K}$-sized TriviaQA, denoted in a dotted vertical line. The violin plots are drawn with randomly chosen $120\mathrm{K}$ samples with $30$ random trials.
@@
Figure 13: Comparison of selective generation methods under a distribution-shift environment with LLaMA3.1-8B-Instruct as a generator ($T=120\mathrm{K},\alpha=0.1$). Here, we consider a single distribution shift from $60\mathrm{K}$-sized NQ to $60\mathrm{K}$-sized TriviaQA, denoted in a dotted vertical line. The violin plots are drawn with randomly chosen $120\mathrm{K}$ samples with $30$ random trials.
@@
Figure 14: Comparison of selective generation methods under a distribution-shift environment with LLaMA3.1-8B-Instruct as a generator ($T=120\mathrm{K},\alpha=0.1$). Here, we consider a single distribution shift from $60\mathrm{K}$-sized TriviaQA to $60\mathrm{K}$-sized NQ, denoted in a dotted vertical line. The violin plots are drawn with randomly chosen $120\mathrm{K}$ samples with $30$ random trials.
F.2.2 Alternating Shift
@@
Figure 15: Comparison of selective generation methods under a distribution-shift environment with GPT-3.5-turbo as a generator ($T=120\mathrm{K},\alpha=0.1$). We consider multiple distribution shifts by alternating $10\mathrm{K}$-sized NQ and TriviaQA 6 times (i.e., NQ, TriviaQA, NQ, TriviaQA and so forth), where shifting points are visualized in dotted vertical lines. The violin plots are drawn with randomly chosen $120\mathrm{K}$ samples with $30$ random trials.
@@
Figure 16: Comparison of selective generation methods under a distribution-shift environment with LLaMA3.1-8B-Instruct as a generator ($T=120\mathrm{K},\alpha=0.1$). We consider multiple distribution shifts by alternating $10\mathrm{K}$-sized NQ and TriviaQA 6 times (i.e., NQ, TriviaQA, NQ, TriviaQA and so forth), where shifting points are visualized in dotted vertical lines. The violin plots are drawn with randomly chosen $120\mathrm{K}$ samples with $30$ random trials.
@@
Figure 17: Comparison of selective generation methods under a distribution-shift environment with LLaMA3.1-8B-Instruct as a generator ($T=120\mathrm{K},\alpha=0.1$). We consider multiple distribution shifts by alternating $10\mathrm{K}$-sized TriviaQA and NQ 6 times (i.e., TriviaQA, NQ, TriviaQA, NQ and so forth), where shifting points are visualized in dotted vertical lines. The violin plots are drawn with randomly chosen $120\mathrm{K}$ samples with $30$ random trials.
F.2.3 Gradual Shift
@@
Figure 18: Comparison of selective generation methods under a distribution-shift environment with GPT-3.5-turbo as a generator ($T=60\mathrm{K},\alpha=0.1$). We consider a distribution shift by gradually shifting from NQ to TriviaQA over time. The violin plots are drawn with randomly chosen $60\mathrm{K}$ samples with $30$ random trials.
@@
Figure 19: Comparison of selective generation methods under a distribution-shift environment with LLaMA3.1-8B-Instruct as a generator ($T=60\mathrm{K},\alpha=0.1$). We consider a distribution shift by gradually shifting from NQ to TriviaQA over time. The violin plots are drawn with randomly chosen $60\mathrm{K}$ samples with $30$ random trials.
@@
Figure 20: Comparison of selective generation methods under a distribution-shift environment with LLaMA3.1-8B-Instruct as a generator ($T=60\mathrm{K},\alpha=0.1$). We consider a distribution shift by gradually shifting from TriviaQA to NQ over time. The violin plots are drawn with randomly chosen $60\mathrm{K}$ samples with $30$ random trials.
F.3 Analysis on Time Horizon and Desired FDR Parameters

### v2

@@ B.4 Adversarial Bandits and Exp3-IX Algorithm
[…]
$\textbf{Reg}_{T}\leq\ell_{\text{max}}\left(2\sqrt{2T|\mathcal{{H}}|\ln|\mathcal{{H}}|}+\left(1+\sqrt{\frac{T|\mathcal{{H}}|}{2\ln|\mathcal{{H}}|}}\right)\ln\delta^{-1}\right).$ (10)
We will exploit Exp3-IX for learning selective generators with partial feedback under an adaptive adversary. See Appendix D for other online learning and bandit algorithms.
B.5 Online Learning and EW Algorithm
Online learning considers to design a learner that adapts to any sequence of full feedback on hypotheses without any or with little assumption on data generation process. In particular, the goal of an online learner is to find a distribution $p_{t}$ over hypotheses with full feedback that minimizes the regret.
Along with many online learning algorithms, we introduce Exponential Weighting (EW) (Littlestone & Warmuth, 1994), which minimizes the regret with finite hypotheses (Algorithm 3). In particular, EW maintains the goodness of weights for each hypothesis in terms of cumulative loss and updates them from feedback. The feedback of a hypothesis $\tau_{t}$ at time $t$ is represented by loss $\ell_{t}(\tau_{t})$ but we can compute this loss for all hypotheses other than a chosen hypothesis $\tau_{t}$ as we directly observe a true label $\mathbbm{1}(G(\mathbf{x}_{t})=_{E}\mathbf{y}_{t})$ at the same time $t$. This EW algorithm satisfies the following regret bound. See Appendix H.1 for a proof.
Theorem 3.
(Littlestone & Warmuth, 1994; Foster & Rakhlin, 2023) Let $\ell_{t}(\cdot)\in[0,\ell_{\text{max}}]$. For any $T\in\mathbb{N}$ and finite hypotheses $\mathcal{{H}}$, Algorithm 3 provides the following expected regret bound if $\eta=\sqrt{{8\ln|\mathcal{{H}}|}/{(\ell_{\text{max}}^{2}T)}}$:
$\sum_{t=1}^{T}\mathbbm{E}_{\tau_{t}\sim p_{t}}\ell_{t}(\tau_{t})-\min_{\tau\in\mathcal{{H}}}\sum_{t=1}^{T}\ell_{t}(\tau)\leq\ell_{\text{max}}\sqrt{{T\ln|\mathcal{{H}}|}/{2}}=\mathcal{{O}}(\ell_{\text{max}}\sqrt{T\ln|\mathcal{{H}}|}),$ (11)
which holds for arbitrary loss sequences, thus, implying that $\mathbbm{E}\textbf{Reg}_{T}$ is also bounded by $\mathcal{{O}}(\ell_{\text{max}}\sqrt{T\ln|\mathcal{{H}}|})$ as the linearity of expectation and Tower property. Moreover, for any $0<\delta<1$, the following regret bound also holds:
$\sum_{t=1}^{T}\ell_{t}(\tau_{t})-\min_{\tau\in\mathcal{{H}}}\sum_{t=1}^{T}\ell_{t}(\tau)\leq\ell_{\text{max}}\sqrt{\frac{T\ln|\mathcal{{H}}|}{2}}+\ell_{\text{max}}\sqrt{\frac{T\ln\delta^{-1}}{2}},$
with probability at least $1-\delta$.
@@ B.5 Online Learning and EW Algorithm
Appendix C Additional Related Work
Constrained Online Learning Several works have studied online learning with constraints. (Mahdavi et al., 2012) first proposes long-term constraints in online convex optimization with time-invariant constraints, which allows to have the constraint function $g_{t}$ to be violated at each time step, i.e., $g(\mathbf{x}_{t})>0$, but eventually satisfied, i.e., $\sum_{t=1}^{T}g(\mathbf{x}_{t})\leq 0$. Some other works solve a similar problem but with time-varying constraints, which is more challenging. (Chen et al., 2017; Sid-Ali et al., 2024) solve a network resource allocation task by viewing this as an constrained optimization problem. Especially, (Sid-Ali et al., 2024) solves the problem without a convex assumption on loss. (Sinha & Vaze, 2024) gives analysis on optimal algorithms for online convex optimization with adversarial constraints. We similarly consider a FDR guarantee to be satisfied over time but with partial feedback.
Conformal Prediction with Bandit Feedback There is little work about selective or conformal prediction in bandit settings. (Ge et al., 2024) proposes an online conformal prediction method with semi-bandit feedback, where we can observe the true label only if it is predicted. Our method follows the practical motivation in having partial feedback idea but we consider selective generation.
Appendix D Additional Method
D.1 Online Selective Generation with Full Feedback
Here, we specifically leverage the conventional exponential weighting (EW) algorithm (Littlestone & Warmuth, 1994) for online learning under adaptive adversaries to re-purpose it for addressing online selective generation with full feedback. This is an oracle method with the performance upper bound of our method with partial feedback. To this end, we provide the reduction of online selective generation with full feedback to the online learning problem. Then, we introduce a modified EW algorithm for selective generation, called EW-SG, followed by its regret bound. Finally, we leverage the proposed conversion lemma, which converts the regret bound of any online learning algorithms into the FDR bound, to show the FDR controllability of EW-SG at a desired level.
D.1.1 Reduction: From Online Selective Generation to Online Learning
@@
models
selective generators $\mathcal{{H}}$
hypotheses $\mathcal{{H}}$
@@
D.1.2 Algorithm and Its Regret Bound
@@ D.2 Regret-to-FDR Conversion
Appendix E Algorithms
Algorithm 2 Procedure for Computing the Loss (4) with Partial Feedback
1:procedure ComputeLoss($\hat{\mathbf{s}},e,\alpha,\lambda$)
2:  $a\leftarrow\mathbbm{1}(\hat{\mathbf{s}}=\texttt{IDK})$
3:  $d\leftarrow\mathbbm{1}(\hat{\mathbf{s}}\neq\texttt{IDK})\cdot e-\alpha\,\mathbbm{1}(\hat{\mathbf{s}}\neq\texttt{IDK})+\alpha$
4:  $\ell\leftarrow\frac{a+\lambda d}{1+\lambda}$
5:  return $\ell$
Algorithm 3 Exponential Weighting (EW) (Littlestone & Warmuth, 1994; Foster & Rakhlin, 2023)
@@ Appendix E Algorithms
Algorithm 4 Exp3 with Implicit eXploration (Exp3-IX) (Neu, 2015)
1:procedure Exp3-IX($T,\mathcal{{H}},\eta_{t},\gamma_{t}$)
@@
7:   Update $w_{t+1}(\tau)\propto\exp(-\eta_{t}\sum_{s=1}^{t}\hat{\ell}_{t}(\tau))$ for all $\tau\in\mathcal{{H}}$
@@
9:   $\ell_{t}\left(\tau,\alpha\right)\leftarrow\frac{a_{t}(\tau)+\lambda d_{t}(\tau,\alpha)}{1+\lambda}$ for all $\tau\in\mathcal{{H}}$
@@
Algorithm 6 Exp3-IX for Online Selective Generation with Partial Feedback (Exp3-IX-SG)
1:procedure Exp3-IX-SG($T,\mathcal{{H}},\alpha,\lambda,\eta,\gamma$)
@@
[…]
Appendix F Experiment Setup
F.1 Computing Environment
@@
F.2 Distribution-shift Environment Setup
@@
Gradual Shift. This setup generates a smooth distribution change over time by drawing individual samples probabilistically between NQ and TriviaQA, with the sampling probability gradually shifting from pure NQ (or TriviaQA) at the start to pure TriviaQA (or NQ) at the end. In particular, given a multinomial distribution $\frac{t}{T},T-\frac{t}{T}]$ for each dataset that changes over time, we first randomly choose a dataset which follows this multinomial distribution, and then we sample one question-answering pair in the sampled dataset. Over the full sequence length, this creates linear interpolation between the two datasets, enabling analysis of model performance under a progressive distribution shift.
F.3 Interactive Environment Setup
@@
To instruct the user-acting agent to generate questions, we use the SQuAD (Rajpurkar et al., 2016) dataset, which consists of many question-answer pairs, each associated with a gold context. Here, we discard the QA pairs and only use $20,962$ context passages, which vary over a range of topics (e.g., science, history, celebrities, entertainment, etc.), forming a diverse set of natural distributions. During the simulation, we randomly select a context from the dataset and provide it to the user-acting agent, which is prompted to continuously generate questions based on the given context. (e.g., Figure 1). To mimic a dynamically shifting environment, e.g., the user’s interest shifts over time, we further instruct the user-acting agent to change the context during the interaction when it determines it is done with the current one (e.g., "When you are done with the current context, say ’SHIFT’ to request a new context.") or after at most 10 question-answering turns. To design this custom workflow, we implement the core functions of the simulation, including memory handling and multi-turn conversation, using LangGraph, a framework built on LangChain (Chase, 2022).
Appendix G Additional Experiments
G.1 Stochastic Environment
@@
Figure 8: Comparison of selective generation methods under a stochastic environment with GPT-3.5-turbo as a generator on TriviaQA ($T=30\mathrm{K},{\alpha=0.08}$). The violin plots are drawn with randomly chosen $30\mathrm{K}$ samples over 100100 random trials.
@@
Figure 9: Comparison of selective generation methods under a stochastic environment with GPT-3.5-turbo as a generator on NQ ($T=30\mathrm{K},{\alpha=0.25}$). The violin plots are drawn with randomly chosen $30\mathrm{K}$ samples with 100100 random trials.
@@
Figure 10: Comparison of selective generation methods under a stochastic environment with LLaMA3.1-8B-Instruct as a generator on NQ ($T=30\mathrm{K},{\alpha=0.25}$). The violin plots are drawn with randomly chosen $30\mathrm{K}$ samples with 100100 random trials.
@@
Figure 11: Comparison of selective generation methods under a stochastic environment with GPT-3.5-turbo as a generator on NQ ($T=30\mathrm{K},{\alpha=0.25}$) and $f_{\texttt{std}}$ as our scoring function. The violin plots are drawn with randomly chosen $30\mathrm{K}$ samples with 100100 random trials.
@@
Figure 12: Comparison of selective generation methods under a stochastic environment with LLaMA3.1-8B-Instruct as a generator on NQ ($T=30\mathrm{K},{\alpha=0.25}$) and $f_{\texttt{std}}$ as our scoring function. The violin plots are drawn with randomly chosen $30\mathrm{K}$ samples with 100100 random trials.
G.2 Distribution-shift Environment
G.2.1 Single Shift
@@
Figure 13: Comparison of selective generation methods under a distribution-shift environment with GPT-3.5-turbo as a generator ($T=30\mathrm{K},\alpha=0.1$). We consider a single distribution shift from NQ to TriviaQA. The violin plots are drawn with randomly chosen $30\mathrm{K}$ samples with 100100 random trials.
@@
Figure 14: Comparison of selective generation methods under a distribution-shift environment with LLaMA3.1-8B-Instruct as a generator ($T=30\mathrm{K},\alpha=0.1$), in a single distribution shift from NQ to TriviaQA. The violin plots are drawn with randomly chosen $30\mathrm{K}$ samples with 100100 random trials.
@@
Figure 15: Comparison of selective generation methods under a distribution-shift environment with LLaMA3.1-8B-Instruct as a generator ($T=30\mathrm{K},\alpha=0.1$), in a single distribution shift from TriviaQA to NQ. The violin plots are drawn with randomly chosen $30\mathrm{K}$ samples with 100100 random trials.
G.2.2 Alternating Shift
@@
Figure 16: Comparison of selective generation methods under a distribution-shift environment with GPT-3.5-turbo as a generator ($T=30\mathrm{K},\alpha=0.1$), in a distribution shift by alternating between NQ and TriviaQA over time, beginning with NQ. The violin plots are drawn with randomly chosen $30\mathrm{K}$ samples with 100100 random trials.
@@
Figure 17: Comparison of selective generation methods under a distribution-shift environment with LLaMA3.1-8B-Instruct as a generator ($T=30\mathrm{K},\alpha=0.1$), in a distribution shift by alternating between NQ and TriviaQA over time, beginning with NQ. The violin plots are drawn with randomly chosen $30\mathrm{K}$ samples with 100100 random trials.
@@
Figure 18: Comparison of selective generation methods under a distribution-shift environment with LLaMA3.1-8B-Instruct as a generator ($T=30\mathrm{K},\alpha=0.1$), in a distribution shift by alternating between NQ and TriviaQA over time, beginning with TriviaQA. The violin plots are drawn with randomly chosen $30\mathrm{K}$ samples with 100100 random trials.
G.2.3 Gradual Shift
(a) FDR over time
(b) Inefficiency over time
(c) FDR distribution
Figure 19: Comparison of selective generation methods under a distribution-shift environment with GPT-3.5-turbo as a generator ($T=30\mathrm{K},\alpha=0.1$), in a gradual distribution shift from NQ to TriviaQA over time. The violin plots are drawn with randomly chosen $30\mathrm{K}$ samples with 100100 random trials.
(a) FDR over time
(b) Inefficiency over time
(c) FDR distribution
Figure 20: Comparison of selective generation methods under a distribution-shift environment with LLaMA3.1-8B-Instruct as a generator ($T=30\mathrm{K},\alpha=0.1$), in a gradual distribution shift from NQ to TriviaQA over time. The violin plots are drawn with randomly chosen $30\mathrm{K}$ samples with 100100 random trials.
(a) FDR over time
(b) Inefficiency over time
(c) FDR distribution
Figure 21: Comparison of selective generation methods under a distribution-shift environment with LLaMA3.1-8B-Instruct as a generator ($T=30\mathrm{K},\alpha=0.1$), in a gradual distribution shift from TriviaQA to NQ over time. The violin plots are drawn with randomly chosen $30\mathrm{K}$ samples with 100100 random trials.
G.3 Analysis on Time Horizon and Desired FDR Parameters